# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@
Abstract
@@
Ergodic quantum many-body systems evolving under unitary time dynamics typically lose memory of their initial state via information scrambling. Here we consider a paradigmatic translationally invariant many-body Hamiltonian of interacting bosons — a Josephson junction array in the transmon regime — in the presence of a strong Floquet drive. Generically, such a time-dependent drive is expected to heat the system to an effectively infinite temperature, featureless state in the late-time limit. However, using numerical exact-diagonalization we find evidence of special ratios of the drive amplitude and frequency where the system develops emergent conservation laws, and approximate integrability. Remarkably, at these same set of points, the Lyapunov exponent associated with the semi-classical dynamics for the coupled many-body equations of motion drops by orders of magnitude, arresting the growth of chaos. We supplement our numerical results with an analytical Floquet-Magnus expansion that includes higher-order corrections, and capture the slow dynamics that controls decay away from exact freezing.
I Introduction
@@
The thermalization dynamics associated with closed quantum many-body systems evolving under their unitary dynamics has been at the center of attention over the past few decades [1, 2, 3, 4]. An intense scrutiny of non-integrable, translationally invariant interacting Hamiltonians, both in theory and experiment, has highlighted the role of the eigenstate thermalization hypothesis (ETH) in the approach to equilibrium. Thermalization proceeds through distinct stages, ranging from local relaxation to scrambling, and is intimately connected to the quantum chaotic nature of the many-body system. Exceptions to such equilibration can arise in strongly disordered problems that can possibly lead to many-body localization (MBL) [5, 6, 7, 8, 9], or in strongly constrained problems that host quantum “scar” states [10, 11, 12]. One of the appeals of arresting the onset of thermalization in many-body systems is their ability to preserve the memory associated with an initial state up to late times. However, the stability of MBL as a distinct phase of matter in the thermodynamic limit remains unclear [13, 14, 15, 16, 17], and the fraction of scar states in the many-body spectrum of many-body Hamiltonians is vanishingly small.
@@
A special class of thermalization dynamics arises in the context of driven many-body systems, such as time-periodic Floquet Hamiltonians. In the absence of strong disorder leading to possible Floquet-MBL or time-crystalline ordered phases [18, 19, 20, 21], such driven Hamiltonians thermalize to a completely featureless state with effectively “infinite-temperature” at asymptotically late times [22, 23]. However, under special circumstances, the heating can be slow with a long regime of a pre-thermal phase [24, 25, 26, 27]. Thus, in generic quantum many-body Hamiltonians with or without a drive, it is difficult to arrest the growth of quantum chaos and prevent thermalization that eventually leads to a complete loss of the memory tied to the initial state.
@@
In parallel developments, based on numerical exact diagonalization (ED) methods and some analytical perturbation theory, it has been suggested that spin models with a bounded local Hilbert space dimension subjected to a strong time-periodic Floquet drive display varying levels of “freezing” at a sequence of special ratios of the drive amplitude and drive frequency  [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. At these points, the system develops approximate emergent conservation laws. The degree to which the system freezes remains unclear, with some conjectures for spin1/212-1/2- 1 / 2 models suggesting an exact freezing [41]. Interestingly, the location of the above freezing-points coincide with the points at which the leading contribution in a Floquet-Magnus expansion vanish. However, there are a number of questions that remain. This includes (i) the role of the local Hilbert space dimension, including an unbounded spectrum, on freezing, (ii) the structure of the higher-order corrections, which controls both the exact location of the freezing point and the residual slow relaxation at freezing, and (iii) the nature of chaos, or lack thereof, at and near freezing. In an accompanying paper [42], some of us have addressed a subset of these questions definitively in an exactly solvable zero-dimensional model of chaotic quantum dots. Furthermore, we also identified a general criterion for freezing, which will be tested in the present manuscript, namely that whenever the commutator between the driving term and the non-integrable part of the static Hamiltonian has a quantized spectrum, the system is prone to display freezing.
Figure 1: Schematic representation of an array of transmons with an unbounded spectrum coupled via capacitive interactions. The external Floquet drive, $f(t)$, couples capacitively to each transmon.
@@
In this manuscript, we address all of these fundamental questions regarding dynamical freezing using a multi-pronged approach. We focus on a paradigmatic model of enormous contemporary interest, namely an array of Josephson junctions (Fig. 1), which forms the fundamental building block of superconducting “transmon” based quantum computing  [43, 44]. A single transmon setup consists of a Josephson junction and a shunting capacitance; the array is constructed out of capacitively coupled transmons. The transmon eigenspectrum is unbounded, but a few of the low lying eigenstates can be chosen to store quantum information. In the transmon regime, the Josephson energy ($E_{J}$) is much greater than the charging energy ($E_{c}$) with typical values in the range $E_{J}/E_{c}\sim 20-80$ [45]. Physically, the transmon array is a quantum version of coupled nonlinear pendula (i.e. oscillators)  [46]. Based on a quantum-to-classical correspondence, it was pointed out that the transmon array shows strong signatures of chaos [46, 47], which can become a future impediment towards scaling up the arrays to tens of thousands of qubits. As the transmons are excited to higher energies, the nonlinearities become more prounounced and trigger a transition from integrable to chaotic motion. In this work we propose that dynamical freezing can be used effectively for suppressing chaos in a large array of transmons, and perhaps other related qubit architectures in the future.
@@
In present day hardware, there are generally two ways by which the parasitic effects of chaos are minimized. The first strategy is to use tunable couplers that control the qubit-qubit interaction by changing the coupler frequency. In this way, qubits are decoupled when off-operation [48]. This approach is adapted by Google’s Sycamore device [49, 50]. A different strategy is to detune the oscillator frequency of neighboring qubits by varying their Josephson energy, which is used by IBM chips [51]. While the first approach stabilizes the system, the additional elements in the circuit add architectural complexity and may introduce new channels of decoherence. In the second approach, the intentional frequency detuning (“disorder”) effectively forces the bosonic quantum many-body system to become many-body localized, and thereby integrable. Whether this is a viable approach with increasing system volume is presently unclear. Regardless of the above protocol, finding new ways to dynamically protect the information stored in individual qubits in the NISQ era devices [52] is a desirable but challenging goal [53].
The remainder of this manuscript is organized as follows. In Sec. II, we introduce the model Hamiltonian of interest and describe the different technical tools that will be used to analyze it in the regime of a strong Floquet drive. These tools include numerical exact diagonalization, a solution for the associated semi-classical Hamilton’s equations of motion, and an analytical Floquet-Magnus expansion. In Sec. III we will discuss the results obtained using exact diagonalization and analyze a variety of diagnostics associated with freezing. We focus on the explicit computation of the semi-classical trajectories and the Lyapunov exponents in Sec. IV. In Sec. V, we extend the computations within a Floquet-Magnus expansion to include the higher-order corrections near freezing and compare against the numerical results. We end in Sec. VI with an outlook towards some open problems. The appendices A and B include a variety of additional technical details.
II Preliminaries: Model and Methods
We introduce the transmon Hamiltonian of interest and discuss the different drive protocols in Sec. II.1, followed by a description of the semi-classical analysis in Sec. II.2. In Sec. II.3, we review key elements of a Floquet-Magnus expansion, which will form the analytical basis for comparing against the numerical results.
II.1 Coupled Transmons
@@
The model of interest to us is an array of coupled transmons [47] in the presence of an external time-periodic drive. The total Hamiltonian is composed of three different parts,
@@
where $\widehat{n}_{i}$ is the Cooper pair number operator, $\widehat{\varphi}_{i}$ is the superconducting phase at site $i$ (which is compact and takes value in the range ($-\pi,\pi]$) , $E_{c}$ is the capacitive charging energy, $E_{J_{i}}$ represent the Josephson energies of the individual transmons (in principle, drawn from some narrow normal distribution), and $V$ is the coupling between nearest-neighbor transmons; see Fig. 1. Physically, $E_{J}$ corresponds to the ability of Cooper pairs to tunnel across the junction. The canonically conjugate operators, $\widehat{n},~{}\widehat{\varphi}$ obey the usual commutation relation $[\widehat{\varphi}_{i},\widehat{n}_{j}]=i\delta_{ij}$, and can be expressed as [54],
@@
where $a_{i}^{\dagger},~{}a_{i}$ represent bosonic creation and annihilation operators, respectively. Note that the spectrum and the associated local Hilbert space for the static Hamiltonian is unbounded from above, which a priori makes the fate of any freezing type phenomenology far from obvious. In particular, the nature of Fock-space (de-)localization, quantum interference, and approach to equilibration via chaos, can be distinctly different in the transmon model compared to spin-models. Interestingly, a Floquet drive-induced many-body localization has been analyzed in a one-dimensional weakly disordered bosonic Hamiltonian [55], where the drive effectively suppresses the static hopping amplitude [56] and pushes the system towards the MBL phase. In contrast, this work focuses on a translationally invariant Hamiltonian, where the static many-body system is fully ergodic and the emergent (approximate) integrability arises via a distinct mechanism.
For simplicity, we have ignored the offset charge(s) [57], $n_{g}(t)$, which are known to be a source of decoherence in charge qubits  [58, 59]. In what follows and unless otherwise stated, we will assume a translationally invariant Hamiltonian with uniform Josephson energies, $E_{J}$; we will briefly comment on the results including the role of randomness in these couplings at the end. We will restrict our attention to the usual transmon regime characterized by $E_{J}\gg E_{c}$ with an eye towards making connection to present day hardware, but the freezing phenomenology is not solely restricted to this hierarchy of energy scales.
@@
Our focus in the remainder of this manuscript will be on the quantum dynamics, chaos and freezing (or lack thereof) for the model introduced in Eq. 1a due to the drive in Eq. 3a-3b as a function of $A/\omega$. We will start by performing an ED analysis in Sec. III below. Given the unbounded local Hilbert-space, ${\cal{H}}_{D}$, we will analyze the eigenspectrum of $H(t)$ with a varying truncation scheme while expanding the $\cos(...)$ non-linearity in powers of $\widehat{\varphi}$; see also Appendix B.
@@
For the numerical results presented in the remainder of this manuscript, we will use the IBM parameters [60], $E_{J}=12.58$  GHz , $E_{c}=330$  MHz, $V=50$  MHz. The transmon regime is defined by $E_{J}\gg E_{c}$, where the oscillator frequency for each transmon is given by $\nu_{i}=\sqrt{8E_{J}E_{c}}-E_{c}$ and is of the order of $\nu\sim 5$ GHz.
II.2 Semi-Classical Limit
The dynamics of the coupled transmons is amenable to a semi-classical analysis, where we replace the non-commuting operators $\widehat{\varphi}$ and $\widehat{n}$ by classical real variables, and treating the commutator bracket as a Poisson bracket, $\{\varphi_{i},n_{j}\}=\delta_{ij}$, respectively. Recall that in the absence of a Floquet drive, the semi-classical dynamics of transmon arrays has been analyzed, and the behavior is generically expected to be chaotic [61, 46] without substantial frequency detuning or tunable coupling [47]. Our goal will be to study how the Floquet drive in the absence of any other strategy helps arrest the growth of chaos, and protect the memory associated with generic initial states.
The Hamilton equations of motion for the coupled set of phase and number variables in Eq. 1a are given by,
where $\sum_{j\in\text{adj}_{i}}$ extends over the nearest neighbors of $i$. In the language of pendula the interaction between the transmons is a momentum-momentum interaction that is due to the capacitive coupling. In addition to obtaining the full dynamics associated with $\varphi_{i}(t),~{}n_{i}(t)$, we will also diagnose chaos by computing the Lyapunov exponent [62]. Specifically, we focus on the maximal Lyapunov exponent (MLE), $\lambda_{L}$, which determines the rate at which trajectories that are initially “close” by diverge with time, $\delta\pi(t)\approx\delta\pi\exp{(\lambda_{L}t)}$. By contrasting the Lyapunov exponent for the static vs. driven transmon arrays, we will highlight in Sec. IV how dynamical freezing effectively arrests the growth of chaos at special values of $A/\omega$.
II.3 Floquet Theory
Our time-dependent Hamiltonian $H(t)=H_{0}+H_{\rm{D}}(t)$ consists of two parts — the static piece, $H_{0}\equiv{\cal{H}}_{\widehat{n}}^{0}+{\cal{H}}_{\widehat{\varphi}}^{0}$, and the driven piece, $H_{\rm{D}}(t)\equiv f(t)\mathcal{H}^{\rm{drive}}_{\widehat{n}}$. Due to the discrete time translational symmetry associated with the period $T\equiv 2\pi/\omega$, the solutions of the Schrödinger equation can be written as |$\ket{\psi_{\alpha}(t)}=e^{-i\epsilon_{\alpha}t}\ket{u_{\alpha}(t)}$, with time-periodic Floquet wavefunction, |$\ket{u_{\alpha}(t+T)}=\ket{u_{\alpha}(t)}$, and Floquet quasi-energy, $\epsilon_{\alpha}\in(-\omega/2,\omega/2]$, respectively. The Floquet-Schrödinger equation is written as [63, 64],
[…]
Figure 2: Stroboscopic time evolution of normalized $\widehat{N}(t)$ for different driving strengths for the square drive in Eq. 3a. The system size is $L=8$, with an on-site Hilbert space dimension $\mathcal{H}_{D}=3$. The drive frequency $\omega=9.32$ GHz, and other parameters are as introduced in the main text. The initial state corresponds to the: (a) ground state, or thermal pure state with (b) $\beta^{-1}=0.79E_{J}$, (c) $\beta^{-1}=1.19E_{J}$, of $H(0)$. At freezing, there is a slow decay of the normalized $\langle\widehat{N}(t)\rangle$ over time (inset).
In the limit of both large driving amplitude and frequency, $H_{\rm eff}$ can be obtained from the modified magnus expansion [33], which we review now. We consider moving into the co-moving frame of the driving term, via the unitary transform
$W(t)=\exp(-i\theta(t)\mathcal{H}_{\widehat{n}}^{\text{drive}})\,,$ (8)
[…]
The Magnus expansion [33] states that the leading order Floquet effective Hamiltonian is given by the one-period average of $\mathcal{H}_{\text{mov}}(t)$:
For the model under consideration, we find
Here $\dots$ denote higher order corrections in $1/\omega$ which will be studied in further detail in Sec. V; see also Appendix A. Therefore, we find at a set of isolated points, the non-commuting $\varphi$-dependent terms vanish, and the effective Hamiltonian becomes seemingly integrable (at leading order):
${\rm{Case~{}I:}}$ $~{}~{}\frac{A}{\omega}\approx 2n,\ {n\in\mathbb{Z}},$ (14a)
Figure 3: Stroboscopic evolution of half-chain entanglement entropy at (red) and away from (blue) freezing for a square-wave drive for three different values of the initial temperature, $\beta^{-1}$. All of the system parameters are same as in Fig. 2. In the inset, we plot the relative change in entropy. The initial states are thermal pure states constructed out of $H(0)$, which includes the drive Hamiltonian.
III Results from Exact Diagonalization
@@
starting from a specific initial state |$\ket{\psi(0)}$, which yields the associated time evolved state, |$\ket{\psi(t)}$. Freezing is identified by the asymptotic late-time behavior of this quantity, which is analyzed as a function of the parameter $A/\omega$. Recall that without the drive, the Hamiltonian $H_{0}\equiv{\cal{H}}_{\widehat{n}}^{0}+{\cal{H}}_{\widehat{\varphi}}^{0}$ is known to be ergodic and chaotic [46], and since $\widehat{N}$ is not a conserved operator its expectation value is expected to relax rapidly. We also compute the time averaged (over Floquet cycles) quantity,
@@
at and away from freezing, which will help determine the shift of the freezing point due to higher order corrections.
We have performed the ED computations for a variety of system sizes, $L(=6,8)$, and on-site Hilbert space dimension, $\mathcal{H}_{D}(=2,3,4)$. We initialize the computation in the time domain with different initial states, |$\ket{\psi(0)}$, which are either eigenstates of $H(0)$ (which includes the drive term), or thermal superposition states constructed out of these, $\propto\sum_{n}e^{-\beta\epsilon_{n}}|n\rangle$. Unless otherwise stated, the default choice for the initial state is the ground state of $H(0)$. We notice that most eigenstates of $H(0)$, except the states in the middle of the spectrum, can be frozen at $A/\omega$ close to the prediction of the leading order Magnus expansion. For many-body states in the middle of the spectrum, the expectation values for $\braket{\widehat{n}_{i}}$ become progressively smaller and the entanglement entropy saturates (see below) in a fashion that makes it numerically difficult to assess the extent of freezing.
@@
where $\rho_{\frac{1}{2}}$ is the density matrix of one half of the chain, computed by tracing out the other half. The entanglement dynamics between different parts of a system serves as a useful diagnostic for identifying the growth of quantum chaos [65]. A non-integrable Floquet system that thermalizes generically at late times, shows a saturating behavior for $S_{\rm ent}$. The saturation value depends on the Hilbert space dimension, and is described by a volume law scaling [66]. However, when the system becomes dynamically frozen, the system develops approximate integrability and leads to extremely slow dynamics. This behavior should be evident in the overall suppression and slow growth of $S_{\rm ent}$. Note that at the freezing point, due to Hilbert space fragmentation [67], the quantum state cannot access the full Hilbert space. Instead, it can only evolve under the invariant Krylov subspace. The dimension of this subspace is much smaller than the dimension of the Hilbert space, and this can lead to Rabi-type oscillations in the entanglement entropy. Finally, we note that the exact nature of the residual dynamics in the frozen state can, in principle, be dependent on the choice of the initial state, which will be analyzed below.
In order to highlight the salient features of the freezing phenomenology that emerges under the square vs. cosine-wave drives, we discuss these separately below in the following two subsections.
III.1 Square-Wave Drive
@@
In Fig. 2, we plot the stroboscopic time evolution for the operator $\widehat{N}$ (defined in Eq. 15), normalized by its initial value. We start either with the ground state of the total Hamiltonian at time $t=0$, or a thermal superposition of the eigenstates of $H(0)$, defined as |$\ket{\Psi(0)}=\sum_{n}\exp{(-\beta\epsilon_{n})}\ket{\psi_{n}}$. When the amplitude of the drive is away from freezing, the system loses memory of the initial state rapidly. On the other hand, at the freezing point, there is an approximate emergent conservation law associated with $\widehat{N}$, and the system retains its memory up to late times. However, we emphasize that this emergent conservation is only approximate, and in Sec. V we calculate the corrections to freezing within a Magnus expansion, which are small in $\mathcal{O}(1/\omega)$. Moreover, these corrections can also result in a shift of the location for freezing. Interestingly, the corrections can depend on the choice of initial states, which therefore also affects the robustness of the associated freezing phenomenology.
@@
In Fig. 3 we show the stroboscopic time evolution of the growth of half chain many-body quantum entanglement entropy ($S_{\rm ent}$) for different intial states. Away from freezing, $S_{\rm ent}$ grows rapidly and saturates to its equilibrated value. At freezing, the system becomes approximately integrable, and the growth of EE becomes slow while its saturation value depends on the Hilbert space dimension of the fragmented sector.
III.2 Oscillatory Drive
For the cosine drive, let us first present results for the numerically time-averaged quantity introduced in Eq. 16 in Fig. 4. We find that the choice of initial state has a strong effect on the location of freezing, i.e. ${\cal{F}}$ peaks at a value of $A/\omega$ that is renormalized away from the value predicted by the zeroth-order Magnus expansion in Eq. 14b. We will revisit this renormalization from the point of view of the higher order corrections in ${\cal{O}}(1/\omega)$ in Sec. V.

### v2

# resub_editorial_supplementary_v2

@@
[…]
resub_editorial_supplementary_v2
Arresting Quantum Chaos Dynamically in Transmon Arrays
Rohit Mukherjee Department of Physics, Cornell University, Ithaca NY 14853. Department of Physics, Indian Institute of Technology, Kanpur 208016, India These authors contributed equally to this work Haoyu Guo Department of Physics, Cornell University, Ithaca NY 14853. These authors contributed equally to this work Keiran Lewellen Department of Physics, Cornell University, Ithaca NY 14853. Debanjan Chowdhury Department of Physics, Cornell University, Ithaca NY 14853. Lead contact Correspondence: debanjanchowdhury@cornell.edu
Summary
@@
Ergodic quantum many-body systems evolving under unitary time dynamics typically lose memory of their initial state via information scrambling. Here we consider a paradigmatic translationally invariant many-body Hamiltonian of interacting bosons — a Josephson junction array in the transmon regime — in the presence of a strong Floquet drive. Generically, such a time-dependent drive is expected to heat the system to an effectively infinite temperature, featureless state in the late-time limit. However, using numerical exact-diagonalization we find evidence of special ratios of the drive amplitude and frequency where the system develops emergent conservation laws, and approximate integrability. Remarkably, at these same set of points, the Lyapunov exponent associated with the semi-classical dynamics for the coupled many-body equations of motion drops by at least an order of magnitude, arresting the growth of chaos. We supplement our numerical results with an analytical Floquet-Magnus expansion that includes higher-order corrections, and capture the slow dynamics that controls decay away from exact freezing.
Keywords
Floquet many-body dynamics, Quantum thermalization, Quantum Chaos, Transmon devices, Floquet-Magnus expansion, Semiclassical dynamics
Introduction
@@
The thermalization dynamics of closed quantum many-body systems has been a major focus in recent decades 1, 2, 3, 4. In non-integrable systems, the eigenstate thermalization hypothesis (ETH) is central to the approach to equilibrium, where thermalization proceeds from local relaxation to scrambling, often driven by quantum chaos. However, thermalization can be interrupted in cases of strong disorder, leading to many-body localization (MBL) 5, 6, 7, 8, 9, 10, or by the presence of quantum “scar” states in constrained systems 11, 12, 13, 14, 15, 16. One of the appeals of arresting the onset of thermalization in many-body systems is their ability to preserve the memory associated with an initial state up to late times. However, the stability of MBL as a distinct phase of matter in the thermodynamic limit remains unclear 17, 18, 19, 20, 21, and the fraction of scar states in the many-body spectrum of many-body Hamiltonians is vanishingly small.
@@
A special class of thermalization dynamics arises in the context of driven many-body systems, such as time-periodic Floquet Hamiltonians. While generically such driven Hamiltonians thermalize to a completely featureless state with effectively “infinite-temperature” at asymptotically late times 22, 23, under special circumstances, the heating can be slow with a long regime of a pre-thermal phase 24, 25, 26, 27.
@@
In an important example of such phenomena, recent studies on strongly driven spin models with bounded local Hilbert spaces show evidence of dynamical “freezing” at a sequence of special ratios of the drive amplitude and drive frequency, leading to approximate emergent conservation laws 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41. Interestingly, the location of the above freezing-points coincide with the points at which the leading contribution in a Floquet-Magnus expansion vanish. However, there are a number of questions that remain. This includes (i) the role of the local Hilbert space dimension, including an unbounded spectrum, on freezing, (ii) the structure of the higher-order corrections, which controls both the exact location of the freezing point and the residual slow relaxation at freezing, and (iii) the nature of chaos, or lack thereof, at and near freezing. In an accompanying paper 42, some of us have addressed a subset of these questions definitively in an exactly solvable zero-dimensional model of chaotic quantum dots. Furthermore, we also identified a general criterion for freezing, which will be tested in the present manuscript.
@@
In this manuscript, we address many of these basic questions regarding dynamical freezing using a multi-pronged approach. We focus on a paradigmatic model of enormous contemporary interest, namely an array of Josephson junctions (Fig. LABEL:fig:Schemetic), which forms the fundamental building block of superconducting “transmon” based quantum computing  43, 44. A single transmon setup consists of a Josephson junction and a shunting capacitance; the array is constructed out of capacitively coupled transmons. The transmon eigenspectrum is unbounded, but a few of the low lying eigenstates can be chosen to store quantum information. Recent work has indicated that such transmon arrays shows strong signatures of chaos 45, 46, 47, 48, 49, which can become a future impediment towards scaling up the arrays to tens of thousands of qubits. As the transmons are excited to higher energies, the nonlinearities become more pronounced and trigger a transition from integrable to chaotic motion. In this work we propose that dynamical freezing can be used effectively for suppressing chaos in a large array of transmons, and perhaps other related qubit architectures in the future.
@@
Presently, chaos is managed by either tunable couplers, as in Google’s Sycamore device 50, 51, or frequency detuning, as in IBM chips 52. While the first approach stabilizes the system, the additional elements in the circuit add architectural complexity and may introduce new channels of decoherence. In the second approach, the intentional frequency detuning (“disorder”) effectively forces the system to become many-body localized, and thereby integrable. Whether this is a viable approach with increasing system volume is presently unclear. Regardless of the above protocol, finding new ways to dynamically protect the information stored in individual qubits in the NISQ era devices 53 is a desirable but challenging goal 54, 55.
Results
Coupled Transmons
@@ Results
The model of interest to us is an array of coupled transmons 46 in the presence of an external time-periodic drive. The total Hamiltonian is composed of three different parts,
@@
where $\widehat{n}_{i}$ is the Cooper pair number operator, $\widehat{\varphi}_{i}$ is the superconducting phase at site $i$ (which is compact and takes value in the range ($-\pi,\pi]$) , $E_{c}$ is the capacitive charging energy, $E_{J_{i}}$ represent the Josephson energies of the individual transmons (in principle, drawn from some narrow normal distribution), and $V$ is the coupling between nearest-neighbor transmons; see Fig. LABEL:fig:Schemetic. Physically, $E_{J}$ corresponds to the ability of Cooper pairs to tunnel across the junction. The canonically conjugate operators, $\widehat{n},~{}\widehat{\varphi}$ obey the usual commutation relation $[\widehat{\varphi}_{i},\widehat{n}_{j}]=i\delta_{ij}$, and can be expressed as 56,
@@
where $a_{i}^{\dagger},~{}a_{i}$ represent bosonic creation and annihilation operators, respectively. Note that the spectrum and the associated local Hilbert space for the static Hamiltonian is unbounded from above, which a priori makes the fate of any freezing type phenomenology far from obvious. In particular, the nature of Fock-space (de-)localization, quantum interference, and approach to equilibration via chaos, can be distinctly different in the transmon model compared to spin-models. Interestingly, a Floquet drive-induced many-body localization has been analyzed in a one-dimensional weakly disordered bosonic Hamiltonian 57, where the drive effectively suppresses the static hopping amplitude 58 and pushes the system towards the MBL phase. In contrast, this work focuses primarily on translationally invariant Hamiltonians, where the static many-body system is fully ergodic and the emergent (approximate) integrability arises via a distinct mechanism.
For simplicity, we have ignored the offset charge(s) 59, $n_{g}(t)$, which are known to be a source of decoherence in charge qubits  60, 61. In what follows and unless otherwise stated, we will assume a translationally invariant Hamiltonian with uniform Josephson energies, $E_{J}$. We will comment on the results of breaking this translational invariance in Sec. LABEL:sec:IBM_simulations. We will restrict our attention to the usual transmon regime characterized by $E_{J}\gg E_{c}$ with an eye towards making connection to present day hardware, but the freezing phenomenology is not solely restricted to this hierarchy of energy scales.
@@
Our focus in the remainder of this manuscript will be on the quantum dynamics, chaos and freezing (or lack thereof) for the model introduced in Eq. (1a) due to the drive in Eq. (3a)-(3b) as a function of $A/\omega$. We will start by presenting results for the exact diagonalization in Sec. LABEL:results:ED below. Given the unbounded local Hilbert-space, ${\cal{H}}_{D}$, we will analyze the eigenspectrum of $H(t)$ with a varying truncation scheme while expanding the $\cos(...)$ non-linearity in powers of $\widehat{\varphi}$; see Supplemental Methods A for details.
@@
For the numerical results presented in the remainder of this manuscript, we will use the IBM parameters 62, $E_{J}/h=12.58$  GHz , $E_{c}/h=330$  MHz, $V/h=50$  MHz. Here to clarify the units we have explicitly restored Planck’s constant $h$, but throughout the remainder of the text we will work in units of $h=1$. The transmon regime is defined by $E_{J}\gg E_{c}$, where the oscillator frequency for each transmon is approximately given by
$h\nu_{i}=\sqrt{8E_{J}E_{c}}$ (4)
and is of the order of $\nu\sim 5$ GHz.
Exact Diagonalization
@@
starting from a specific initial state |$\ket{\psi(0)}$, which yields the associated time evolved state, |$\ket{\psi(t)}$. Freezing is identified by the asymptotic late-time behavior of this quantity, which is analyzed as a function of the parameter $A/\omega$. Recall that without the drive, the Hamiltonian $H_{0}\equiv{\cal{H}}_{\widehat{n}}^{0}+{\cal{H}}_{\widehat{\varphi}}^{0}$ is known to be ergodic and chaotic 45, and since $\widehat{N}$ is not a conserved operator its expectation value is expected to relax rapidly. We also compute the time averaged (over Floquet cycles) quantity,
@@
at and away from freezing, which serves as a proxy for measuring the emergent conservation of $\hat{N}$ and will help determine the shift of the freezing point due to higher order corrections. Here delimited-⟨$\langle$..$\rangle$ … ⟩ represents the expectation value with respect to some chosen state and the “$\overline{...}$” corresponds to a time average over many Floquet-cycles.
In the main text, we focus on a one-dimensional system of 8 transmons with a local Hilbert space dimension of 3 for our exact-diagonalization computations. However, we have thoroughly examined the freezing phenomenon by varying both the system size and the onsite Hilbert space dimension in Supplemental Methods A. Our findings indicate that the freezing point remains highly robust as a function of the varying system size and Hilbert space dimension.
We initialize the computation in the time domain with different initial states, |$\ket{\psi(0)}$, which are either eigenstates of $H(0)$ (which includes the drive term), or thermal superposition states constructed out of these, $\propto\sum_{n}e^{-\beta\epsilon_{n}}|n\rangle$. Unless otherwise stated, the default choice for the initial state is the ground state of $H(0)$. We notice that most eigenstates of $H(0)$, except the states in the middle of the spectrum, can be frozen at $A/\omega$ close to the prediction of the leading order Magnus expansion. For many-body states in the middle of the spectrum, the expectation values for $\braket{\widehat{n}_{i}}$ become progressively smaller and the entanglement entropy saturates (see below) in a fashion that makes it numerically difficult to assess the extent of freezing. The ground state or the low temperature thermal superposition states of $H(0)$ have relatively large expectation values for $\braket{\widehat{n}_{i}}$, so by choosing to initialize in these state, we manage to avoid the numerical issues stated above.
@@
where $\rho_{\frac{1}{2}}$ is the density matrix of one half of the chain, computed by tracing out the other half. The entanglement dynamics between different parts of a system serves as a useful diagnostic for identifying the growth of quantum chaos 63. A non-integrable Floquet system that thermalizes generically at late times, shows a saturating behavior for $S_{\rm ent}$. The saturation value depends on the Hilbert space dimension, and is described by a volume law scaling 64. However, when the system becomes dynamically frozen, the system develops approximate integrability and leads to extremely slow dynamics. This behavior should be evident in the overall suppression and slow growth of $S_{\rm ent}$. Finally, we note that the exact nature of the residual dynamics in the frozen state can, in principle, be dependent on the choice of the initial state, which will be analyzed below.
In order to highlight the salient features of the freezing phenomenology that emerges under the square vs. cosine-wave drives, we discuss these separately below in the following two subsections. We will see that, while freezing occurs for both drives, it is quantitatively stronger in the case of the square-wave drive. This behavior aligns well with what our Magnus-expansion based theory predicts and stems from the unique form of the Fourier coefficients of the square-wave drive. (See Supplemental Methods B for a detailed derivation.)
Square-Wave Drive
@@
In Fig. LABEL:fig:ndtstrobo(a),(b),(c), we plot the stroboscopic time evolution for the operator $\widehat{N}$ (defined in Eq. (5)), normalized by its initial value. We start either with the ground state of the total Hamiltonian at time $t=0$, or a thermal superposition of the eigenstates of $H(0)$, defined as |$\ket{\Psi(0)}=\sum_{n}\exp{(-\beta\epsilon_{n})}\ket{\psi_{n}}$. When the amplitude of the drive is away from freezing, the system loses memory of the initial state rapidly. On the other hand, at the freezing point, there is an approximate emergent conservation law associated with $\widehat{N}$, and the system retains its memory up to late times. However, we emphasize that this emergent conservation is only approximate, and in Sec. LABEL:results:FM we calculate the corrections to freezing within a Magnus expansion, which are small in $\mathcal{O}(1/\omega)$. Moreover, these corrections can also result in a shift of the location for freezing which is more pronounced in the cosine wave driving, that we will revisit using the Magnus expansion. Interestingly, the corrections can depend on the choice of initial states, which therefore also affects the robustness of the associated freezing phenomenology.
@@
In Fig. LABEL:fig:ndtstrobo(d),(e),(f) we show the stroboscopic time evolution of the growth of half chain many-body quantum entanglement entropy ($S_{\rm ent}$) for different initial states. Away from freezing, $S_{\rm ent}$ grows rapidly and saturates to its equilibrated value. At freezing, the system becomes approximately integrable, and the growth of EE becomes slow while its saturation value depends on the Hilbert space dimension. In the inset of Fig. LABEL:fig:ndtstrobo(d),(e),(f) we plot the relative change in entropy. With increasing temperature, which incorporates a superposition over a larger number of states from the many-body spectrum, the relative change increases at freezing but remains much smaller than away from freezing. Additionally, for various initial states away from freezing, the saturation value remains nearly the same, primarily dictated by the total Hilbert space dimension.
Cosine-Wave Drive
For the cosine drive, let us first present results for the emergent conservation of $N$ in Fig. LABEL:fig:cosonefidelity, where we plot $\mathcal{F}$ (defined in Eq. (6)) for different choice of initial states. We find several salient features, which we now discuss one by one. First, compared to the square-wave drive case, we observe a sizeable renormalization of the location of the freezing point (marked by black arrows in Fig. LABEL:fig:cosonefidelity(a)) compared to the prediction of the zeroth-order Magnus expansion (marked by dashed lines in Fig. LABEL:fig:cosonefidelity(a)). In Sec. LABEL:results:FM, we will show that such renormalizations of the freezing point are controlled by the corrections in the Magnus expansion, which also depends on the choice of initial state.
Second, the states in the middle of the spectrum are harder to freeze (see e.g. the less pronounced peak for the pink curve in Fig. LABEL:fig:cosonefidelity(a)). There are two possible explanations for the reduced extent of freezing: (i) These initial states have a higher entanglement entropy (as shown later in Fig. LABEL:fig:cosinendt) and hence longer range of entanglement. Therefore, the state is more susceptible to higher order corrections in the Magnus expansion, which consists of longer operator strings. (ii) These states have a smaller value of $\braket{\hat{N}}$, and is therefore more susceptible to fluctuations of $\hat{N}$.
Third, apart from the freezing peaks predicted by the leading-order Magnus expansion, we also observed secondary peaks that are located in between the freezing points. These peaks can also be explained within the Magnus framework. At these secondary locations, half of the non-integrable terms in the Hamiltonian are effectively switched off. A more detailed description of the scenario is given in Supplemental Methods C.